OpenStudy (anonymous):

a parking lot has a rectangular area of 40,000yd^2. The length is 200 yd more than twice the width. What are the dimensions of the lot? So far I have width=x length=2x+200 so A=lw so 40,000=(2x+200)(x)

OpenStudy (anonymous):

then i took 2x^2 +200x=40,000 everything after that is a mess

OpenStudy (anonymous):

answer should be 100x400 how do i get there?

OpenStudy (anonymous):

Set it equal to 0: \[2x^{2} + 200x - 40000 = 0\] Then factor: \[(2x - 200)(x+200)\]

OpenStudy (anonymous):

\[x=100\]\[(x+200)(x-100)=0\]\[x^2+100x-20000=0\]